NEXT
Make New Entry,
Make Followup Entry
User name jmammei
Log entry time 10:24:13 on March 02, 2011
Entry number 349614
keyword=RC daily summary
Run Coordinator Daily Summary
Available Beam in Use:
Day - 0.98 hours
Swing - 3.93 hours
Owl - 7.03 hours
Lost time Problems
- Day - Mostly BNA because of beam studies, recover from cryo 2K cold
box trip
- Swing - Troubleshooting DT issue, recover from cryo 2K cold box trip
- Owl - BeO and Al target changes, reboot DAQ
Issues
- Cryo trip
delayed our return after beam studies (waiting for R Dipole) on day/swing
shifts
- Dead Time high at kinematic setting #2 - 20% with 800Hz (see run
plan)
- Timing
issues discovered while troubleshooting DT
- Production overnight
Shift Summaries
ABUs
A copy of this log entry has been emailed to: doug, sparveri, danez, sarty@ap.stmarys.ca